# 2024_subroc.bib

### v1

@@ A.1 Average Ranking Loss
It follows that the following expression implies Equation 18.
@@
It remains to show that Equation 20 holds.
It is trivial that the average of a multiset $[m_{1},\dots,m_{|M|}]=M$ of real numbers $m_{i}\in\mathbb{R}$ is not greater than the maximum of $M$:
@@
Equation 20 follows from Equation 21 and Equation 22, concluding the proof for Equation 16.
Proof for Equation 17
@@
Let $i=\arg\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}$ and $C^{\prime}=(C\circledast\{(y,\hat{y})\in C\mid y=0\wedge\hat{y}\geq\hat{y}_{i}\})\cup[c_{i}]$. $C^{\prime}$ consists of the positive instance $c_{i}$ and exactly the elements of $C$ that define the non-zero summands in the definition of $PEN_{i}$ (see main text). By definition of the ranking loss, $C^{\prime}$ demonstrates the existence of a sub-multiset of $C$ for which Equation 23 holds.
This concludes the proof for Equation 17. ∎
@@
Lemma 10 states that a tight soft classifier performance measure upper bound for the (loss-type) average ranking loss is given by the function
@@
It follows from Lemma 7 that a tight optimistic estimate for $\varphi^{asl}$ is given by
$oe_{\varphi^{asl}}(p)=\max_{i\in\{1,\dots,|sg(p)|\}}\{RL_{i}(sg(p)^{\{y,\hat{y}\}})\}$
@@
$c=-ARL(\mathcal{I}^{\{y,\hat{y}\}})$ is constant with respect to the fixed dataset $D$. It follows from Lemma 9 that a tight optimistic estimate for $\varphi^{rasl}$ is given by
$oe_{\varphi^{rasl}}(p)=\max_{i\in\{1,\dots,|sg(p)|\}}\{RL_{i}(sg(p)^{\{y,\hat{y}\}})\}-ARL(\mathcal{I}^{\{y,\hat{y}\}})$
@@ A.2 Area Under the ROC Curve
holds. This condition also holds for any sub-multiset of $C$, that contains at least one instance with $y_{i}=1$ and one with $y_{i}=0$ each (such that ROC AUC is defined), because of the universal quantifier. Let $C^{\prime}$ be such a sub-multiset. We continue with a proof by contradiction, showing that the premise $ROCAUC(C^{\prime})<\frac{1}{2}$ implies that Equation 24 is not true. Firstly the diagonal ROC curve with only supporting points $(0,0)$ and $(1,1)$ has an AUC of $\frac{1}{2}$. Clearly any ROC curve with only supporting points $(0,0)$, $(1,1)$ and additional points above or on the diagonal ROC curve has an AUC of $\frac{1}{2}+x$ with $x\in\mathbb{R}_{\geq 0}$. Therefore $ROCAUC(C^{\prime})<\frac{1}{2}$ implies that at least one supporting point of the ROC curve of $C^{\prime}$ lies below the diagonal ROC curve. This is equivalently expressed by
@@
$\Leftrightarrow$ $\exists t\in\{\hat{y}\mid(y,\hat{y})\in C^{\prime}\}:$ $\frac{1}{P(C^{\prime})}\cdot|[(y,\hat{y})$ $\in C^{\prime}\mid y=1\wedge\hat{y}>t]|<$
$\frac{1}{N(C^{\prime})}\cdot|[(y,\hat{y})$ $\in C^{\prime}\mid y=0\wedge\hat{y}>t]|.$
@@
If there exist instances $(y,\hat{y})$ and $(y^{\prime},\hat{y}^{\prime})$ in $C^{\prime}$ that fulfill this expression, then they also fulfill the following expressions:
$\Rightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $y>y^{\prime}\wedge\hat{y}<\hat{y}^{\prime}$
$\Rightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $(y<y^{\prime}\wedge\hat{y}>\hat{y}^{\prime})\vee(y>y^{\prime}\wedge\hat{y}<\hat{y}^{\prime})$
The negation of Equation 24 is equivalent:
$\Leftrightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $\neg(y\geq y^{\prime}\wedge\hat{y}\leq\hat{y}^{\prime})\vee\neg(y\leq y^{\prime}\wedge\hat{y}\geq\hat{y}^{\prime})$
$\Leftrightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $\neg((y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime}))$
$\Leftrightarrow$ $\neg\forall(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ ::: $(y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime})$
Since in case $b_{ROCAUC}(C)=\frac{1}{2}$ we know that Equation 24 holds, we can conclude that $ROCAUC(C^{\prime\prime})\geq\frac{1}{2}$ holds for any sub-multiset of $C$, that contains at least one instance with $y_{i}=1$ and one with $y_{i}=0$ each (such that ROC AUC is defined). Therefore $b_{ROCAUC}$ is a soft classifier performance measure lower bound for ROC AUC in this case.
The case $b_{ROCAUC}(C)=0$ is present if and only if $C$ contains erroneous predictions, meaning that there exists at least two instances $(y,\hat{y})$ and $(y^{\prime},\hat{y}^{\prime})$ in $C$ such that $y=0$, $y^{\prime}=1$ and $\hat{y}>\hat{y}^{\prime}$. The sub-multiset of $C$ that only contains these two instances (once each) has an ROC AUC of 00. This is the overall minimum of the ROC AUC. $b_{ROCAUC}$ consequently meets the criteria of a tight soft classifier performance measure lower bound for $ROCAUC$ in case $b_{ROCAUC}(C)=0$. ∎
@@
Lemma 2 states that a tight soft classifier performance measure lower bound for the (score-type) area under the ROC curve is given by $b_{ROCAUC}$.
It follows from Lemma 8 that a tight optimistic estimate for $\varphi^{ROCAUC}$ is given by
$oe_{\varphi^{ROCAUC}}(p)=-b_{ROCAUC}(sg(p)^{\{y,\hat{y}\}})$
@@
$c=ROCAUC(\mathcal{I}^{\{y,\hat{y}\}})$ is constant with respect to the fixed dataset $D$. It follows from Lemma 9 that a tight optimistic estimate for $\varphi^{rROCAUC}$ is given by
$oe_{\varphi^{rROCAUC}}=ROCAUC(\mathcal{I}^{\{y,\hat{y}\}})-b_{ROCAUC}(sg(p)^{\{y,\hat{y}\}})$
@@ A.3 Area Under the Precision-Recall Curve
$T(C)$ :=$:=\{(-\infty,\hat{y}_{1})\}\cup\{(\hat{y}_{i-1},\hat{y}_{i})\mid i\in\{2,\dots,|C|\}\wedge\hat{y}_{i-1}\neq\hat{y}_{i}\}$
$T_{<\hat{y}_{minpos}}(C)$ :=$:=\{(t_{lo},t_{hi})\in T(C)\mid t_{hi}<\hat{y}_{minpos}\}$
$T_{\geq\hat{y}_{minpos}}(C)$ :=$:=\{(t_{lo},t_{hi})\in T(C)\mid t_{hi}\geq\hat{y}_{minpos}\}$
@@
In particular, Equation 30 also holds for any sub-multiset $C^{\prime}$ of $C$ with corresponding $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$. Therefore
@@
=$=\left\{(0,1),(0,0),\left(1,\frac{1}{1+N(C_{worst})}\right)\right\}.$
@@
$\mathfrak{P}_{worst}^{\prime}$ =$=\left\{(0,0),\left(1,\frac{P(C_{worst})}{P(C_{worst})+N(C_{worst})}\right)\right\}$
=$=\left\{(0,0),\left(1,\frac{1}{1+N(C_{worst})}\right)\right\}.$
Clearly the AUC between the points $(0,1)$ and $(0,0)$ is 00, so
@@
$\mathfrak{P}_{worst}^{\prime}$ are the supporting points of the minimum PR curve described in [boyd_unachievable_2012]. As $\mathfrak{P}_{worst}^{\prime}$ shows, the minimum PR curve only depends on the positive class ratio. By definition of $C_{worst}$ there is no $C^{\prime}\subseteq C_{\hat{y}\geq\hat{y}_{minpos}}$ such that the positive class ratio of $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$ is strictly lower than that of $C_{worst}$. Also no point on the linear interpolation of $\mathfrak{P}_{worst}^{\prime}$ is strictly above the correct PR interpolation of $\mathfrak{P}_{worst}^{\prime}$ as shown in the following.
The points on the correct interpolation between the two points in $\mathfrak{P}_{worst}^{\prime}$ according to [boyd_unachievable_2012] are
@@
Lemma 4 states that a tight soft classifier performance measure lower bound for the (score-type) area under the PR curve is given by $b_{PRAUC}$.
It follows from Lemma 8 that a tight optimistic estimate for $\varphi^{PRAUC}$ is given by
$oe_{\varphi^{PRAUC}}(p)=-b_{PRAUC}(sg(p)^{\{y,\hat{y}\}})$
@@
$c=PRAUC(\mathcal{I}^{\{y,\hat{y}\}})$ is constant with respect to the fixed dataset $D$. It follows from Lemma 9 that a tight optimistic estimate for $\varphi^{rPRAUC}$ is given by
$oe_{\varphi^{rPRAUC}}=PRAUC(\mathcal{I}^{\{y,\hat{y}\}})-b_{PRAUC}(sg(p)^{\{y,\hat{y}\}})$
@@
$\Leftrightarrow$ 11 11 $\geq\left(1-\frac{1}{P(C)}\right)^{\frac{\beta}{\alpha}}+\frac{1}{P(C)+N(C)}$ (32)
@@
11 11 $\geq\left(1-\frac{1}{P(C)}\right)^{1}+\frac{1}{P(C)+N(C)}$
@@
Therefore Equation 32 holding in case $\alpha=\beta$ implies that it also holds in case $\alpha\leq\beta$. We conclude that the weighting $w$ does not decrease when $c$ is removed from $C$.
@@
Assume $N(C)=0\vee P(C)=0$ (Equal to the case in the $cb$ definition returning 00).
@@
$b_{w}$ is a tight upper bound in this case since it also returns 00:
@@ B.2 Quality Parameters of Results Across Multiple Datasets
Table 4: Overview of result set properties for combinations of datasets and performance measures for an xgboost classifier. (Continuation of Table 1 in the main text)
Mean Exceptionality
@@
Top-5/ Significant/ Original
Dataset Metric     b f     b f     b f     b f     b f
Mushroom n=8124 NCR=0.52 ROC AUC     0.00 -    203 -    0.76 -    0.28 -    0/5 0/0/100
PR AUC     0.00 0.00    35 57    0.17 0.49    0.14 -    0/5 1/1/100
@@
Dataset Interestingness Measure
@@ B.3 Case Study: Adult
The following are full result sets of the searches on the Adult dataset, that are also reported in Table 1 in the main text in terms of aggregated statistics. Note in particular the amount of overlap between patterns in the result sets from the base SubROC setting compared to the subgroups in the results of the full SubROC framework. Moreover values of the main text table are reflected in the extreme performance results, cover sizes and negative class ratios.
@@
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
@@
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
@@
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
@@
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
Table 10: Result set of full SubROC framework with ROC AUC.
@@
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482
@@
[…]
$0$emptyset 477.4517 0.9355 0.8298 11305 0.2482

### v2

@@ A.1 Average Ranking Loss
It follows that the following expression implies Equation˜18.
@@
It remains to show that Equation˜20 holds.
It is trivial that the average of a multiset $m_{1},\dots,m_{|M|}]=M$ of real numbers $m_{i}\in\mathbb{R}$ is not greater than the maximum of $M$:
@@
Equation˜20 follows from Equation˜21 and Equation˜22, concluding the proof for Equation˜16.
Proof for Equation˜17
@@
Let $i=\arg\max_{i\in\{1,\dots,|C|\}}\{RL_{i}(C)\}$ and $^{\prime}=(C\circledast\{(y,\hat{y})\in C\mid y=0\wedge\hat{y}\geq\hat{y}_{i}\})\cup[c_{i}]. CC^{\prime}$ consists of the positive instance $c_{i}$ and exactly the elements of $C$ that define the non-zero summands in the definition of $PEN_{i}$ (see main text). By definition of the ranking loss, $C^{\prime}$ demonstrates the existence of a sub-multiset of $C$ for which Equation˜23 holds.
This concludes the proof for Equation˜17. ∎
@@
Lemma˜10 states that a tight soft classifier performance measure upper bound for the (loss-type) average ranking loss is given by the function
@@
It follows from Lemma˜7 that a tight optimistic estimate for $\varphi^{asl}$ is given by
[…]
@@
[…]
@@ A.2 Area Under the ROC Curve
holds. This condition also holds for any sub-multiset of $C$, that contains at least one instance with $y_{i}=1$ and one with $y_{i}=0$ each (such that ROC AUC is defined), because of the universal quantifier. Let $C^{\prime}$ be such a sub-multiset. We continue with a proof by contradiction, showing that the premise $ROCAUC(C^{\prime})<\frac{1}{2}$ implies that Equation˜24 is not true. Firstly the diagonal ROC curve with only supporting points $(0,0)$ and $(1,1)$ has an AUC of $\frac{1}{2}$. Clearly any ROC curve with only supporting points $(0,0)$, $(1,1)$ and additional points above or on the diagonal ROC curve has an AUC of $\frac{1}{2}+x$ with $x\in\mathbb{R}_{\geq 0}$. Therefore $ROCAUC(C^{\prime})<\frac{1}{2}$ implies that at least one supporting point of the ROC curve of $C^{\prime}$ lies below the diagonal ROC curve. This is equivalently expressed by
@@
[…]
@@
If there exist instances (y,y^)($\hat{y}$) and $(y^{\prime},\hat{y}^{\prime})$ in $C^{\prime}$ that fulfill this expression, then they also fulfill the following expressions:
$\Rightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ $:$ $y>y^{\prime}\wedge\hat{y}<\hat{y}^{\prime}$
$\Rightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ $:$ $(y<y^{\prime}\wedge\hat{y}>\hat{y}^{\prime})\vee(y>y^{\prime}\wedge\hat{y}<\hat{y}^{\prime})$
The negation of Equation˜24 is equivalent:
$\Leftrightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ $:$ $\neg(y\geq y^{\prime}\wedge\hat{y}\leq\hat{y}^{\prime})\vee\neg(y\leq y^{\prime}\wedge\hat{y}\geq\hat{y}^{\prime})$
$\Leftrightarrow$ $\exists(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ $:$ $\neg((y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime}))$
$\Leftrightarrow$ $\neg\forall(y,\hat{y}),(y^{\prime},\hat{y}^{\prime})\in C^{\prime}$ $:$ $(y<y^{\prime}\rightarrow\hat{y}\leq\hat{y}^{\prime})\wedge(y>y^{\prime}\rightarrow\hat{y}\geq\hat{y}^{\prime})$
Since in case $b_{ROCAUC}(C)=\frac{1}{2}$ we know that Equation˜24 holds, we can conclude that $ROCAUC(C^{\prime\prime})\geq\frac{1}{2}$ holds for any sub-multiset of $C$, that contains at least one instance with $y_{i}=1$ and one with $y_{i}=0$ each (such that ROC AUC is defined). Therefore $b_{ROCAUC}$ is a soft classifier performance measure lower bound for ROC AUC in this case.
The case $b_{ROCAUC}(C)=0$ is present if and only if $C$ contains erroneous predictions, meaning that there exists at least two instances (y,y^)($\hat{y}$) and $(y^{\prime},\hat{y}^{\prime})$ in $C$ such that $y=0$, $y^{\prime}=1$ and $\hat{y}>\hat{y}^{\prime}$. The sub-multiset of $C$ that only contains these two instances (once each) has an ROC AUC of 0. This is the overall minimum of the ROC AUC. $b_{ROCAUC}$ consequently meets the criteria of a tight soft classifier performance measure lower bound for $ROCAUC$ in case $b_{ROCAUC}(C)=0$. ∎
@@
Lemma˜2 states that a tight soft classifier performance measure lower bound for the (score-type) area under the ROC curve is given by $b_{ROCAUC}$.
It follows from Lemma˜8 that a tight optimistic estimate for $\varphi^{ROCAUC}$ is given by
[…]
@@
[…]
@@ A.3 Area Under the Precision-Recall Curve
$T(C)$ $:=\{(-\infty,\hat{y}_{1})\}\cup\{(\hat{y}_{i-1},\hat{y}_{i})\mid i\in\{2,\dots,|C|\}\wedge\hat{y}_{i-1}\neq\hat{y}_{i}\}$
$T_{<\hat{y}_{minpos}}(C)$ $:=\{(t_{lo},t_{hi})\in T(C)\mid t_{hi}<\hat{y}_{minpos}\}$
$T_{\geq\hat{y}_{minpos}}(C)$ $:=\{(t_{lo},t_{hi})\in T(C)\mid t_{hi}\geq\hat{y}_{minpos}\}$
@@
In particular, Equation˜30 also holds for any sub-multiset $C^{\prime}$ of $C$ with corresponding $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$. Therefore
@@
$=\left\{(0,1),(0,0),\left(1,\frac{1}{1+N(C_{worst})}\right)\right\}.$
@@
$\mathfrak{P}_{worst}^{\prime}$ $=\left\{(0,0),\left(1,\frac{P(C_{worst})}{P(C_{worst})+N(C_{worst})}\right)\right\}$
$=\left\{(0,0),\left(1,\frac{1}{1+N(C_{worst})}\right)\right\}.$
Clearly the AUC between the points $(0,1)$ and $(0,0)$ is 0, so
@@
$\mathfrak{P}_{worst}^{\prime}$ are the supporting points of the minimum PR curve described in [16]. As $\mathfrak{P}_{worst}^{\prime}$ shows, the minimum PR curve only depends on the positive class ratio. By definition of $C_{worst}$ there is no $C^{\prime}\subseteq C_{\hat{y}\geq\hat{y}_{minpos}}$ such that the positive class ratio of $C_{\hat{y}\geq\hat{y}_{minpos}}^{\prime}$ is strictly lower than that of $C_{worst}$. Also no point on the linear interpolation of $\mathfrak{P}_{worst}^{\prime}$ is strictly above the correct PR interpolation of $\mathfrak{P}_{worst}^{\prime}$ as shown in the following.
The points on the correct interpolation between the two points in $\mathfrak{P}_{worst}^{\prime}$ according to [16] are
@@
Lemma˜4 states that a tight soft classifier performance measure lower bound for the (score-type) area under the PR curve is given by $b_{PRAUC}$.
It follows from Lemma˜8 that a tight optimistic estimate for $\varphi^{PRAUC}$ is given by
[…]
@@
[…]
@@
$\Leftrightarrow$ $1$ $\geq\left(1-\frac{1}{P(C)}\right)^{\frac{\beta}{\alpha}}+\frac{1}{P(C)+N(C)}$ (32)
@@
$1$ $\geq\left(1-\frac{1}{P(C)}\right)^{1}+\frac{1}{P(C)+N(C)}$
@@
Therefore Equation˜32 holding in case $\alpha=\beta$ implies that it also holds in case $\alpha\leq\beta$. We conclude that the weighting $w$ does not decrease when $c$ is removed from $C$.
@@
Assume $N(C)=0\vee P(C)=0$ (Equal to the case in the $cb$ definition returning 0).
@@
$b_{w}$ is a tight upper bound in this case since it also returns 0:
@@ B.2 Quality Parameters of Results Across Multiple Datasets
Table 5: Overview of result set properties for combinations of datasets and performance measures for an xgboost classifier. (Continuation of Table˜1 in the main text)
Mean Exception- ality
@@
Filtered-5/ Significant/ Top-100
Dataset
Metric
b
f
b
f
b
f
b
f
b
f
Mushroom n=8124 NCR=0.52 ROC     0.00 -    203 -    0.76 -    0.28 -    0/5 0/0/100
PR     0.00 0.00    35 57    0.17 0.49    0.14 -    0/5 1/1/100
@@
Dataset
Interestingness
Measure
@@ B.3 Case Study: Adult
The following are full result sets of the searches on the Adult dataset, that are also reported in Table˜1 in the main text in terms of aggregated statistics. Note in particular the amount of overlap between patterns in the result sets from the base SubROC setting compared to the subgroups in the results of the full SubROC framework. Moreover values of the main text table are reflected in the extreme performance results, cover sizes and negative class ratios.
@@
159.2829
$\text{capital-gain}\in[0.0,114.0)\>\wedge$
$\text{capital-loss}\in[0.0,213.0)\>\wedge$
$\text{marital-status}=\text{Married-civ-spouse}$
372.3026
0.8409
0.7780
4309
0.3892
157.6086
$\text{capital-gain}\in[0.0,114.0)\>\wedge$
$\text{marital-status}=\text{Married-civ-spouse}$
370.6283
0.8473
0.8053
4628
0.4114
142.0611
$\text{capital-loss}\in[0.0,213.0)\>\wedge$
$\text{marital-status}=\text{Married-civ-spouse}$
355.0808
0.8611
0.8346
4961
0.4328
141.2313
$\text{marital-status}=\text{Married-civ-spouse}$
354.2510
0.8660
0.8508
5280
0.4496
138.0929
$\text{capital-gain}\in[0.0,114.0)\>\wedge$
$\text{capital-loss}\in[0.0,213.0)\>\wedge$
$\text{marital-status}=\text{Married-civ-spouse}\>\wedge$
$\text{native-country}=\text{United-States}$
351.1126
0.8275
0.7748
3894
0.4004
0.0000
$\emptyset$
213.0197
0.9728
0.9241
11305
0.2482
@@
0.9241
$\text{education}=\text{Some-college}\>\wedge$
$\text{fnlwgt}\in[107160.0,132053.0)\>\wedge$
$\text{marital-status}=\text{Divorced}\>\wedge$
$\text{workclass}=\text{Private}$
1.0000
0.8439
0.0000
21
0.0476
0.9241
$\text{education}=\text{Some-college}\>\wedge$
$\text{fnlwgt}\in[178811.0,196338.0)\>\wedge$
$\text{relationship}=\text{Not-in-family}\>\wedge$
$\text{sex}=\text{Male}$
1.0000
0.8857
0.0000
32
0.0312
0.9241
$\text{capital-gain}\in[0.0,114.0)\>\wedge$
$\text{education}=\text{Some-college}\>\wedge$
$\text{fnlwgt}\in[107160.0,132053.0)\>\wedge$
$\text{marital-status}=\text{Divorced}$
1.0000
0.8770
0.0000
29
0.0345
0.9241
$\text{education-num}\in[10,11)\>\wedge$
$\text{fnlwgt}\in[107160.0,132053.0)\>\wedge$
$\text{marital-status}=\text{Divorced}\>\wedge$
$\text{race}=\text{White}$
1.0000
0.8703
0.0000
27
0.0370
0.9241
$\text{education-num}\in[10,11)\>\wedge$
$\text{fnlwgt}\in[107160.0,132053.0)\>\wedge$
$\text{marital-status}=\text{Divorced}\>\wedge$
$\text{workclass}=\text{Private}$
1.0000
0.8439
0.0000
21
0.0476
0.0000
$\emptyset$
213.0197
0.9728
0.9241
11305
0.2482
@@
0.9395
$\text{education}=\text{Bachelors}\>\wedge$
$\text{hours-per-week}\in[44,50)\>\wedge$
$\text{occupation}=\text{Prof-specialty}\>\wedge$
$\text{relationship}=\text{Husband}$
14.0000
0.0333
0.4400
26
0.9615
0.9395
$\text{education-num}\in[13,14)\>\wedge$
$\text{hours-per-week}\in[44,50)\>\wedge$
$\text{occupation}=\text{Prof-specialty}\>\wedge$
$\text{relationship}=\text{Husband}$
14.0000
0.0333
0.4400
26
0.9615
0.9371
$\text{education-num}\in[14,15)\>\wedge$
$\text{hours-per-week}\in[50,55)\>\wedge$
$\text{occupation}=\text{Exec-managerial}\>\wedge$
$\text{sex}=\text{Male}$
13.0000
0.0357
0.5357
29
0.9655
0.9371
$\text{education}=\text{Masters}\>\wedge$
$\text{hours-per-week}\in[50,55)\>\wedge$
$\text{occupation}=\text{Exec-managerial}\>\wedge$
$\text{sex}=\text{Male}$
13.0000
0.0357
0.5357
29
0.9655
0.9311
$\text{education}=\text{Masters}\>\wedge$
$\text{hours-per-week}\in[50,55)\>\wedge$
$\text{marital-status}=\text{Married-civ-spouse}\>\wedge$
$\text{occupation}=\text{Exec-managerial}$
11.0000
0.0417
0.5600
26
0.9615
0.0000
$\emptyset$
213.0197
0.9728
0.9241
11305
0.2482
@@
609186.2021
$\text{marital-status}=\text{Married-civ-spouse}$
354.2510
0.8660
0.8508
5280
0.4496
389481.0168
$\text{relationship}=\text{Husband}$
315.8451
0.8619
0.8485
4638
0.4495
34597.0291
$\text{sex}=\text{Male}$
223.1345
0.9541
0.9053
7581
0.3109
9394.5568
$\text{capital-gain}\in[0.0,114.0)$
216.3884
0.9711
0.9015
10331
0.2126
4270.2607
$\text{capital-gain}\in[0.0,114.0)\>\wedge$
$\text{sex}=\text{Male}$
228.3650
0.9505
0.8764
6821
0.2708
0.0000
$\emptyset$
213.0197
0.9728
0.9241
11305
0.2482
Table 11: Result set of full SubROC framework with ROC AUC. (Extended version of Table˜2 in the main text)
@@
316.1949
$\text{marital-status}=\text{Married-civ-spouse}$
354.2510
0.8660
0.8508
5280
0.4496
286.2383
$\text{relationship}=\text{Husband}$
315.8451
0.8619
0.8485
4638
0.4495
67.9166
$\text{capital-gain}\in[0.0,114.0)\>\wedge$
$\text{marital-status}=\text{Married-civ-spouse}$
370.6283
0.8473
0.8053
4628
0.4114
64.1438
$\text{sex}=\text{Male}$
223.1345
0.9541
0.9053
7581
0.3109
63.0874
$\text{capital-gain}\in[0.0,114.0)$
216.3884
0.9711
0.9015
10331
0.2126
0.0000
$\emptyset$
213.0197
0.9728
0.9241
11305
0.2482
@@
460.8622
$\text{marital-status}=\text{Married-civ-spouse}$
354.2510
0.8660
0.8508
5280
0.4496
420.2511
$\text{relationship}=\text{Husband}$
315.8451
0.8619
0.8485
4638
0.4495
87.2251
$\text{occupation}=\text{Exec-managerial}$
71.7942
0.9121
0.9027
1518
0.4862
79.5800
$\text{occupation}=\text{Prof-specialty}$
69.6873
0.9104
0.8978
1467
0.4649
74.2654
$\text{hours-per-week}\in[50,55)$
58.9793
0.8939
0.8844
1113
0.4582
0.0000
$\emptyset$
213.0197
0.9728
0.9241
11305
0.2482